If 87 people attend a concert and tickets for adults cost $3.75 while tickets for children cost $3.25 and total receipts for the concert was $298.75, how many of each went to the concert?
Answer by TimothyLamb(4379)

x = adults
y = kids
---
x + y = 87
3.75x + 3.25y = 298.75
---
put the system of linear equations into standard form
---
x + y = 87
3.75x + 3.25y = 298.75
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = adults = 32
y = kids = 55
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---